A basketball player applies a force F = 65 lb to the rim at A. Determine the equivalent force couple system at point B, which is at the center of the rim mounting bracket on the backboard


1
Expert's answer
2021-10-24T18:20:21-0400


from 5-12-13 triangle:

"F_z=12F/13=12\cdot65/13=60" lb

"F_{xy}=5F/13=5\cdot65/13=25" lb


"F_x=F_{xy}sin60\degree=25sin60\degree=21.65" lb

"F_y=F_{xy}cos60\degree=25cos60\degree=12.5" lb


Since there is only one force acting on the rim, the resultant force will be equal to that force "\vec{F}" , and the couple will be equal to the moment of "\vec{F}" about point B:

"\vec{R}=\vec{F}=21.56i-12.5j-60k" lb

"M_B=\vec{r}_{BA}\times \vec{F}"


coordinates of points A and B:

"A(-9sin30\degree,9cos30\degree,0)=A(-4.5,7.79,0)"

"B(0.-15,-4)"


"\vec{r}_{BA}=-4.5i+22.75j+4k"


"M_B=\begin{vmatrix}
 i & j&k \\
 -4.5 & 22.79&4\\
21.56&-12.5&-60
\end{vmatrix}="


"=(-22.79\cdot60+4\cdot12.5)i-(4.5\cdot60-4\cdot21.56)j+(4.5\cdot12.5-22.79\cdot21.56)k="

"=-1317.4i-183.76j-435.10k" lb"\cdot"in